a=-4, b=2, c=-3
1)5b-(a+c)
2)100(2b+a)
Please explain every step.
thank you

Answer by Flannery(124)

Substitute numbers for variables:
5(2)-(-4+(-3))=
Work from the inner most parentheses outward:
(-4+(-3)) is the same as -4-3 because the addition of a negative number is a subtraction. So (-4+(-3)) is (-4-3). Now we have:
5(2)-(-4-3)=
Distribute the implied negative one. Multiplying two negative numbers always results in a positive answer. So -1(-4-3) is 4+3. We now have:
5(2)+4+3=
Finish removing parentheses by multiplying five and two:
10+4+3=
10+7=
17
Same again:
100(2times2+(-4))= one hundred times (two times two plus (negative four))
Innermost parentheses first. Adding the negative four means subtracting four.
100(2times2-4))=
100(4-4)=
finish ding what's in the parentheses:
100(0)=
0
The trick is to do your parentheses in order from innermost to outermost and to know when to change the signs.
I change signs first if needed, then distribute if needed,and then do remaining operations within the parentheses.
These answers were checked by Microsoft Mathematics.